OpenStudy (anonymous):

Solve for X algebraically: 3^(2X) + 3^(X+1) - 4 = 0

OpenStudy (anonymous):

This involves logarithms because you are solving for the exponent in both of them to solve for x.

OpenStudy (anonymous):

First we need to rewrite equation

OpenStudy (anonymous):

\[(3^x)^2 + 3^x*3-4=0\]

OpenStudy (anonymous):

Are you getting this ???

OpenStudy (anonymous):

Hi, thank you for the response. Does the +1 in the exponent become *3 ?

OpenStudy (anonymous):

You could have solved this with logarithms.

OpenStudy (anonymous):

yes correct Now we need to substitute 3^x = m ( any variable ) so can you try to rewrite equation ??

OpenStudy (anonymous):

Like a U substitution with U= 3^X ?

OpenStudy (anonymous):

How you got 12 ? we just need to replace 3^x by U okay ??

OpenStudy (anonymous):

U^2 + U *3 -4 =0 ??? ??????

OpenStudy (anonymous):

Perfect Now we have to solve this quadratic equation can you try ?

OpenStudy (anonymous):

It isn't in standard form. I don't know what to do with the part with: *3 -4

OpenStudy (anonymous):

you can use quadratic formula for it can you try ?

OpenStudy (anonymous):

Hello .... ????

OpenStudy (anonymous):

what would be C ?

OpenStudy (anonymous):

X would be equal to 0.

OpenStudy (anonymous):

x = –1, 3.